# https://vision.ist.i.kyoto-u.ac.jp/

### v2

@@ 6.2 LSP
6.2.1 Implementation Details.
@@
As shown in Tab. 6, LSP outperforms the other two methods in most cases. Particularly, the accuracy of MonoCInIS drops on Cityscapes since it overfits the training dataset, i.e., KITTI and cannot address the domain gap. In contrast, thanks to the large training data and various augmentations, our LSP obtains the highest generalizability and robustly estimates dimensions on any dataset. Learning not only car height but also width and length also improves the prediction accuracy as this facilitates to understand overall car sizes.
Table 6: Generalizability on KITTI and Cityscapes. We evaluate LSP on Absolute Relative Error for each dimension (height/width/length), comparing with the fixed car height prior (1.59 m, following [20]) and MonoCInIS [19] trained on KITTI. LSP robustly surpass other methods in most cases and can predict accurately regardless of the datasets.
@@
KITTI
Fixed Height [20] 0.065 / - / - 0.064 / - / - 0.065 / - / - 0.065 / - / -
@@
Cityscapes
Fixed Height [20] 0.093 / - / - 0.093 / - / - 0.076 / - / - 0.092 / - / -
@@ 7 Conclusion
In this work, we proposed StableCamH, a novel training framework to make any MDE model learn metric scale. StableCamH aggregates the scale information obtained from a learned object size prior into the camera height and uses it as pseudo-supervision formulated as optimization across frames and epochs by leveraging its invariance. We showed that StableCamH enables monocular depth estimators to grasp metric scale regardless of network architectures and achieves state-of-the-art absolute-scale depth estimation accuracy on both the KITTI and the Cityscapes dataset, and demonstrated that it can leverage any dataset for training. We also proposed LSP, which can be easily trained on a large-scale dataset without manual annotations and robustly estimates vehicle dimensions from their appearances. We showed that the learned size prior is superior to a fixed one and outperforms a similar 3D object detection model. StableCamH can be leveraged as a drop-in training scheme for any monocular depth estimation method. We hope it will be a versatile tool for road scene understanding research.
@@ 0.A.1 Dataset Preparation
We use the images from the front camera in Argoverse2 [46], Lyft [21], A2D2 [13], and DDAD [16] as training datasets. As mentioned in Sec. 6.1 in the main paper since the focal lengths are essential for metric MDE, we resize the image and align the focal lengths to 1000 px. In order to facilitate the training process, we crop the images to 832$\times$512 resolution. These resizing and cropping are also conducted during evaluation on KITTI Eigen split [8].
To make the training easier and more efficient, we eliminate the frames where the camera is stationary, since training with these frames greatly degrades the inference performance [15]. We achieve this elimination by looking at pairs of successive frames and simply checking the number of pixels with large differences in normalized intensities. Images with a small number of pixels with large intensity changes are deemed stationary and discarded from the training dataset. Since Argoverse2 is captured at a higher frame rate than the other datasets, we treat every other frame as successive frames.
For optimization during training, StableCamH requires a sufficient number of images captured with the same camera height. We select such sequences that include at least 2,500 frames. At every end of the training epoch, we optimize each camera height individually. In total, we use 235,341 images for training.
@@ 0.A.2 Training Settings
We train Monodepth2 [15] with ResNet50 on StableCamH with the same batch size and initial learning rate as used during training on the KITTI dataset (Sec. 6.1 in the main paper). Considering the gap of training samples between KITTI and the mixed datasets, we set $\tau_{\mathrm{mid}}=4$ and train the model for 8 epochs while reducing the learning rate by half every 3 epochs.
Appendix 0.B Ablation of LSP
@@
As specified in Sec. 5 in the main paper, during training of LSP, we perform beam/instance occlusion and truncation augmentations illustrated in  Fig. 3 along with common color jittering and blur. The beam occlusion simulates objects occluded by poles including traffic lights and trees. Their segmentation mask is divided into more than two regions. The instance occlusion augmentation is for objects occluded by the other stuffs including vehicles and people. The truncation one is for objects that are partially cropped out of images. In  Tab. 7 we ablate the effectiveness of these augmentations. The results show that each augmentation contributes to improving the prediction accuracy both on the KITTI and Cityscapes datasets. This indicates that they enhance the robustness to any input image.
@@
Appendix 0.C Additional Quantitative Results on KITTI
In the main paper, we trained the multiple networks with StableCamH on the KITTI dataset [12] and evaluated them on the KITTI Eigen split [8] which is the most widely used, comparing the weakly-supervised methods. We also evaluate the same models with the improved ground-truth depth maps for KITTI introduced in [37]. These improved depth maps are composed of 652 out of 697 test frames in the Eigen split. The results shown in Tab. 8 demonstrate that the models trained with StableCamH still surpass weakly-supervised methods, both with and without the median scaling, as well as when evaluating on the KITTI Eigen split. It should be emphasized that VADepth [48] trained with StableCamH (row 9 and 18) outperforms the same network supervised with the measured camera height (row 4 and 13). This proves that the pseudo camera height obtained via the camera height optimization is more precise than the measured one.
Table 8: Evaluation on KITTI improved ground truth [37] (640$\times$192 resolution). The models trained with StableCamH outperform weakly-supervised methods in both results, with and without the median scaling [59].
@@ Appendix 0.C Additional Quantitative Results on KITTI
[…]
@@
Figure 4: Additional qualitative comparison on KITTI. We compare estimated depth maps of Monodepth2 [15] trained with StableCamH (Ours) with the ones of weakly-supervised methods: G2S [3], DynaDepth [56], and VADepth [48]. In error maps, the larger depth errors are represented in red, and smaller ones are depicted in blue. The results show that StableCamH can predict more accurate depth maps in various scenes, compared with the weakly-supervised methods.
Appendix 0.D Additional Qualitative Results on KITTI
We provide additional results of StableCamH in diverse scenes on the KITTI dataset. As shown in Fig. 4, StableCamH can predict more accurate depth maps than the weakly-supervised methods, especially in road and object regions.

### v3

@@ 6.2 LSP
We provide implementation details of LSP in the supplementary material.
@@
As shown in Tab. 6, LSP outperforms the other two methods in most cases. Particularly, the accuracy of MonoCInIS drops on Cityscapes since it overfits to the training dataset, i.e., KITTI and cannot address the domain gap. In contrast, thanks to the large training data and various augmentations, our LSP obtains the highest generalizability and robustly estimates dimensions on any dataset. Learning not only car height but also width and length also improves the prediction accuracy as this facilitates estimation of the overall car sizes.
Table 6: Generalizability on KITTI and Cityscapes. We evaluate LSP on Absolute Relative Error for each dimension (height/width/length), comparing with the fixed car height prior (1.59 m, following [23]) and MonoCInIS [22] trained on KITTI. LSP robustly surpasses other methods in most cases and can predict accurately regardless of the datasets.
@@
KITTI Fixed Height [23] 0.065 / - / - 0.064 / - / - 0.065 / - / - 0.065 / - / -
@@
Cityscapes Fixed Height [23] 0.093 / - / - 0.093 / - / - 0.076 / - / - 0.092 / - / -
@@ 7 Conclusion
In this work, we proposed FUMET, a novel training framework for any MDE model to learn metric scale. FUMET aggregates the scale information obtained from a learned car size prior into the camera height and uses it as pseudo-supervision formulated as optimization across frames and epochs by leveraging its invariance. We showed that FUMET enables monocular depth estimators to embody metric scale regardless of the underlying network architectures and achieves state-of-the-art absolute-scale depth estimation accuracy on both the KITTI and the Cityscapes datasets, and demonstrated that it can leverage any dataset for training. We also proposed LSP, which can be easily trained on a large-scale dataset without manual annotations and robustly estimates vehicle dimensions from their appearances. We showed that the learned size prior is superior to a fixed one and outperforms a similar 3D object detection model. FUMET can be leveraged as a drop-in training scheme for any monocular depth estimation method. We hope it will be a versatile tool for road scene understanding research.
@@ 0.A.1 Dataset Preparation
We use the images from the front camera in Argoverse2 [50], Lyft [24], A2D2 [16], and DDAD [19] for training. As mentioned in Sec. 6.1 since the focal lengths are essential for metric MDE, we resize the image and align the focal lengths to 1000 px. In order to facilitate the training process, we crop the images to 832$\times$512 resolution. These resizing and cropping are also done for evaluation on the KITTI Eigen split [11].
To make the training easier and more efficient, we eliminate frames where the camera is stationary since training with these frames greatly degrades the inference performance [18]. We achieve this elimination by looking at pairs of successive frames and simply checking the number of pixels with large differences in normalized intensities. Images with a small number of pixels with large intensity changes are deemed stationary and discarded from the training dataset. Since Argoverse2 is captured at a higher frame rate than the other datasets, we treat every other frame as successive frames.
For optimization during training, FUMET requires a sufficient number of images captured with the same camera height. We select such sequences with at least 2,500 frames. At the end of each training epoch, we optimize each camera height individually. In total, we use 235,341 images for training.
@@ 0.A.2 Training Settings
We train Monodepth2 [18] with ResNet50 on FUMET with the same batch size and initial learning rate as used during training on the KITTI dataset (Sec. 6.1). Considering the gap of training samples between KITTI and the mixed dataset, we set $\tau_{\mathrm{mid}}=4$ and train for 8 epochs while reducing the learning rate by half every 3 epochs. For training on the mixed dataset and KITTI, we train for 8 epochs with $\tau_{\mathrm{mid}}=3$ and reduce the learning rate by half every 2 epochs.
Appendix 0.B Camera Intrinsics Recovered with COLMAP
Similar to previous self-supervised MDE methods, FUMET requires camera intrinsics to compute the photometric error in training. In the following section, we demonstrate that FUMET successfully trains a model with intrinsics recovered with COLMAP [37] on the KITTI dataset [15]. In Sec. 0.B.2, we also verify that FUMET can leverage in-the-wild videos for training by using COLMAP.
0.B.1 Training on KITTI with Recovered Camera Intrinsics
We run COLMAP with the pinhole camera assumptions on KITTI sequences containing 500 to 800 frames and use the median of the computed intrinsics for training. We train Monodepth2 [18] (ResNet50) on FUMET under the same training settings as described in Sec. 6.1. The results on KITTI (Tab. 7) show that thanks to the accuracy of COLMAP, the model trained with the recovered intrinsics achieves comparable accuracy to that trained with the ground-truth values.
Table 7: Comparison of FUMET trained with ground-truth camera intrinsics versus the one trained with recovered values with COLMAP. Each model is trained and evaluated on KITTI [15, 11] (640$\times$192 resolution). FUMET can train a model with recovered camera intrinsics, achieving accuracy comparable to that trained with ground-truth ones.
Error ($\downarrow$) Accuracy ($\uparrow$)
Intrinsics AbsRel SqRel RMSE RMSElog $\delta\!<\!1.25$ $\delta\!<\!1.25^{2}$ $\delta\!<\!1.25^{3}$
GT 0.108 0.785 4.736 0.195 0.871 0.958 0.981
COLMAP 0.110 0.774 4.721 0.197 0.867 0.958 0.981
0.B.2 Training with In-the-wild Videos
We download two YouTube videos [1, 2] (73,100 frames) and run COLMAP to recover the intrinsics. Since these videos are too long to use all the frames for COLMAP, we manually extract consecutive 300 frames from uncrowded scenes and run COLMAP individually with the pinhole camera assumptions. We train Monodepth2 with FUMET on the dataset composed of the mixed dataset (Argoverse2 [50], Lyft [24], A2D2 [16], and DDAD [19]), KITTI and the YouTube videos. The batch size and the initial learning rate are the same as those used in Sec. 6.1. For the same reason as described in Sec. 0.A.2, we train for 6 epochs with $\tau_{\mathrm{mid}}=2$, reducing the learning rate by half every 2 epochs.
The results on KITTI shown in Tab. 8 demonstrate that adding in-the-wild videos to the training dataset increases the dataset diversity and improves the depth accuracy. This suggests the significance of being able to train without sensor measurements.
Table 8: FUMET can leverage YouTube videos for training and lead to slightly improved accuracy. The results are evaluated on KITTI [11] (832$\times$512 resolution). In the Dataset column, M, K, and Y represent the mixed dataset, KITTI, and YouTube, respectively.
Error ($\downarrow$) Accuracy ($\uparrow$)
Dataset AbsRel SqRel RMSE RMSElog $\delta\!<\!1.25$ $\delta\!<\!1.25^{2}$ $\delta\!<\!1.25^{3}$
M $+$ K 0.082 0.611 4.307 0.149 0.923 0.976 0.989
M $+$ K $+$ Y 0.084 0.608 4.264 0.148 0.923 0.977 0.989
Appendix 0.C Assumption of Vehicle Presence
In contrast to inference time, FUMET leverages observed vehicles and creates pseudo scale supervision for training. However, it does not create pseudo supervision for each frame independently but instead aggregates the scale information into the camera height and optimizes it across the sequence. This allows FUMET to create scale supervision from frames observing vehicles and provide scale supervisory signals even for frames without observed vehicles. In fact, as shown in Fig. 3, although the training dataset of KITTI contains many frames where vehicles are not observed or are scarce, FUMET can train with high depth accuracy.
Figure 3: The proportion of frames with the number of observed cars in the training dataset of KITTI [15]. There is no or only one observed car in nearly half of the frames in the dataset,
Appendix 0.D Comparison to a Supervised Method
Our key contribution is a self-supervision cue for metric depth reconstruction. As such, direct comparison with supervised methods (e.g., ZoeDepth [4]) would be unfair. That said, we can compare with supervised methods trained on our mixed dataset using LiDAR depth. We use ZoeDepth as a representative method for this comparison. We select the architecture most similar to our default model (ResNet50) from its official implementation in terms of the number of parameters and the architecture type, i.e., ConvNet, and skip pre-training performed in the original paper for fairness.
The results on KITTI (Tab. 9) show that FUMET slightly outperforms ZoeDepth. This suggests that it can achieve accuracy comparable to supervised methods. This is likely because the richer self-supervision from the reconstruction loss (Eq. 3) forces the model to recognize scene contexts effectively, more so than sparse LiDAR depth, and enables metric scale inference from them without overfitting.
Table 9: Comparison with ZoeDepth [4] trained on the mixed dataset with LiDAR depth. The results are evaluated on KITTI [11] (832$\times$512 resolution). FUMET slightly outperforms ZoeDepth in terms of depth accuracy even though FUMET does not use ground-truth depth maps.
Error ($\downarrow$) Accuracy ($\uparrow$)
Method Supervision AbsRel SqRel RMSE RMSElog $\delta\!\!<\!\!1.25$ $\delta\!\!<\!\!1.25^{2}$ $\delta\!\!<\!\!1.25^{3}$
ZoeDepth [4] GT 0.120 0.743 5.118 0.182 0.870 0.964 0.988
FUMET - 0.113 0.916 5.009 0.181 0.883 0.961 0.983
Appendix 0.E Implementation Details of LSP
Appendix 0.F Ablation Studies of LSP
@@
As specified in Sec. 5 in the main paper, during training of LSP, we perform beam/instance occlusion and truncation augmentations illustrated in  Fig. 4 along with common color jittering and blur. The beam occlusion simulates objects occluded by poles including traffic lights and trees. Their segmentation mask is divided into more than two regions. The instance occlusion augmentation is for objects occluded by the other things including vehicles and people. Truncation augmentation is for objects that are partially cropped out of images. In  Tab. 10, we ablate the effectiveness of these augmentations. The results show that each augmentation contributes to improving the prediction accuracy both on the KITTI and the Cityscapes datasets, i.e., they enhance the robustness to any input image.
@@
Appendix 0.G Additional Quantitative Results on KITTI
In the main paper, we trained the multiple networks with FUMET on the KITTI dataset [15] and evaluated them on the KITTI Eigen split [11] which is the most widely used, comparing the weakly-supervised methods. We also evaluate the same models with the improved ground-truth depth maps for KITTI introduced in [41]. These improved depth maps are composed of 652 out of 697 test frames in the Eigen split. The results shown in Tab. 11 demonstrate that the models trained with FUMET still surpass weakly-supervised methods, both with and without the median scaling, as well as when evaluating on the KITTI Eigen split. It should be emphasized that VADepth [52] trained with FUMET (row 9 and 18) outperforms the same network supervised with the measured camera height (row 4 and 13). This proves that the pseudo camera height obtained via the camera height optimization is more precise than the measured one.
Table 11: Evaluation on KITTI improved ground truth [41] (640$\times$192 resolution). The models trained with FUMET outperform weakly-supervised methods in both results, with and without the median scaling [63].
@@ Appendix 0.G Additional Quantitative Results on KITTI
Method Supervision Scaling AbsRel SqRel RMSE RMSElog $\delta\!\!<\!\!1.25$ $\delta\!\!<\!\!1.25^{2}$ $\delta\!\!<\!\!1.25^{3}$
@@
Appendix 0.H Additional Qualitative Results on KITTI
We provide additional results of FUMET in diverse scenes on the KITTI dataset. As shown in Fig. 5, FUMET can predict more accurate depth maps than the weakly-supervised methods, especially in road and object regions.
@@ Appendix 0.H Additional Qualitative Results on KITTI
Figure 5: Additional qualitative comparison on KITTI. We compare estimated depth maps of Monodepth2 [18] trained with FUMET to the ones of weakly-supervised methods: G2S [6], DynaDepth [60], and VADepth [52]. In error maps, the larger depth errors are represented in red, and smaller ones are depicted in blue. The results show that FUMET can predict more accurate depth maps in various scenes, compared with the weakly-supervised methods.